LHC 2013 - latest news

Week 3: Machine Coordinators: Gianluigi Arduini, Barbara Holzer

Goal of the week: Setting-up for proton-lead physics

p-Pb commissioning plan (John Jowett)

Date

Summary

Link to report from morning meeting

Mon 21st January
  • injecting 338 bunches/beam fill 3479
  • 8:13 dump during cogging
08:30 meeting
Ion summary (J.Jowett)
Sun 20th January
  •  2:50 - 10:30 3 fills for off-momentum loss maps and asynch dump
  • 10:50 - 12:45 asynch. dump test and B2 loss maps at injection
  • 15:08 stable beam p,Pb fill 3474, 1.5 hours; 13 bunches/beam
  • change to 'ADJUST': fixed and re-calibrated BSRA; verified that the abort gap cleaning is working
  • Dumped beam 2 in order to do loss maps with B1 only: B1H, B1V and negative off-momentum loss maps in collision
  • 19:00 start set-up of p and Pb trains
  • 21:00 trains in the LHC 96 p and 120 Pb bunches
  • 21:22 dump by SIS during ramp due to missing BPM data
09:00 meeting
Sat 19th January  
  • 01:30 Started roman pot alignment in collision
  • 02:40. Beams dumped  by wrong step in ALFA roman pot alignment
  • 03:51 Start injection No ion beam from SPS. Continue with B1
  • 04:30-07:30 Loss maps at injection with B1
  • 06:00 Ion beam available but problem with Beam Dump B2 ==> expert called. Access required to replace PLC.
  • 08:40 access
  • 09:30 Access finished
  • 11:17 12:50 RF problem with ion injection (re-phasing), due to a disabled timing
  • 13:22 - 19:17 fill for RP set-up followed by betatron loss maps and off-momentum loss maps at collision with RP in
  • dry-run and correction of sequence for RP to physics settings
  • 21:45 - 1:49 fill with betatron loss maps at flat top, squeezed and separated and in physics condition with RPs in
09:00 meeting
Fri 18th January
  • 01:45-05:30 Injection of protons and ions, capture, steering
  • 05:30-07:20 Cogging, squeeze
  • 07:20-09:30 go in collision
  • 09:20 Collisions in all IPs
  • 09:30 10:24: Orbit optimization, FBCT re-phasing, collimator alignment check for TCSG.6L7.B2
  • 10:28-14:50. Trip of RSF1.A56B2 during pre-cycle, invesitgations and access to repair.
  • 15:30-17:30 Filling
  • 18:30 Cogging at flat-top completed.
  • 18:30 - 19:20 Loss maps at flat-top (betatronic)
  • 19:20-19:40 Squeeze
  • 19:40-20:46 TCT alignment
  • 20:46-23:00 Loss maps and some alignment verifications
  • 23:00 Colliding in all points
  • 23:50-01:25 TCT alignment and loss maps in collision
08:30 meeting
Thu 17th January
  • 06:00 cryo back start pre-cycle
  • 07:20 Injection
  • Adjust beam in the injectors and set-up oorbit acquisition
  • 12:30 End of ramp
  • 15:00 First set of loss maps at flat-top completed
  • 16:00 Collimator alignment in point 7 completed
  • 17:30 Second set of loss maps completed
  • 17:47 Squeeze completed in one go
  • 18:00-24:00 IR2 aperture measurement
08:30 meeting
A. Suraci / cryo
Wed 16th January
  • 00:03  TDI OK after several movements by piquet (reason fro problem unclear)
  • 03:36 No way to fix problem with Beam Imminent Warning sirens. Manual trigger for the night.
  • 04:13 B2 injection
  • 04:13 lost cryo conditions in sector 34/45 (ciommunication problem)
  • 06:30 Communication problem identified (faulty network switch). IT piquet called.
  • 09:30 IT piquet accessing. BIW siren problem being investigated by expert.
  • 11:30 Second access required for network switch
  • 12:30 Cryo communication problem in point 4 fixed
  • 13:17 Trip of the cold compressor in point 4.
  • afternoon / evening: problems reconnecting cryo in Pt4
08:30 meeting
F. Duval on Power glitch
Tue 15th January
  • 4:12 off-momentum optics at beta*=0.8m
  • 08:28 Power glitch due to a short on 18 kV cable in Meyrin. Al sectors tripped. Cryo in P18 affected.
  • Accesses during the day (CV, ALICE, CMS, ventilation doors)
  • 20:30 Problem with sirenes in point 5 and 6. Not possible to close the machine. Expert called.
  • 21:00 Cryogenics recovered
  • Problem with TDI movement. Expert called.
  • 23:30 Undulator trip (RU.L4). Loss of cryo conditions in the matching section L4.
08:30 meeting
R.Versteegen
Mon 14th January
  • 02:20 beam loss due to instabilities
  • 03:15 new injection
  • 04:25 beta sqeeze, 80cm reached, optics checks
  • 09:20 fixing issues at injection
  • 13:10 set-up injection for 200ns trains of protons
  • 16:40 - 18:20 access for ALICE (and BGI in parallel)
  • preparing injection of Pb trains
  • 20:00 - 22:20 no beam from SPS
  • 22:20 optics with experiment bumps on-momentum during squeeze and collision beam process
08:30 meeting
F.Duval